# Safe Reinforcement Learning In Black-Box Environments Via Adaptive Shielding

### v2

@@ 4.2 ADVICE Execution and Adaptation
Upon completing the shield construction phase, ADVICE can be used as a post-shield within the RL loop. The constructed shield encourages safe environment exploration, guarding the agent throughout its interaction with the environment by ensuring the execution of safe actions.
[…]
@@
At the end of each episode, ADVICE updates its cautiousness level by considering the recent safety performance of the RL agent (line 22). This unique ADVICE characteristic enables moving beyond the static definition of the safety threshold $K$ by automatically adjusting its value based on its recent safe/unsafe behaviour. Accordingly, ADVICE becomes adaptive and allows the agent to explore more when exhibiting safe behaviour for a period of time while being more cautious, thus interfering more, when the RL agent behaves increasingly unsafely.
To assess the agent’s performance over time, ADVICE employs a double sliding window, commonly used in the field of anomaly detection (Tu et al., 2019; Wang et al., 2024). By comparing the current trend of safety violations against a broader historical view, ADVICE can discern whether the current trend deviates from the pattern seen historically. This analysis informs ADVICE to strengthen or relax its cautiousness level during the execution of an episode. The adaptation function is given by:
$K=\begin{cases}\min(K\!+\!1,K^{max}),&\text{if }\mu_{r}\!\!>\!\!\mu_{d}\!+\!\sigma_{d}\\ \max(K\!-\!1,\lceil K^{max}/2\rceil),&\text{if }\mu_{r}\!\!<\!\!\mu_{d}-\sigma_{d}\\ K,&\text{otherwise}\end{cases}$ (4)
@@
If the $\mu_{r}$ value exceeds the sum of $\mu_{d}+\sigma_{d}$, this signifies that the agent incurs safety violations more often than expected. Consequently, $K$ is automatically incremented to adopt a more cautious position. Inversely, if $\mu_{r}$ is less than $\mu_{d}-\sigma_{d}$, it is an indication that the agent acts more cautious than expected. As a result, $K$ is automatically reduced to allow the agent to explore more freely. The safety threshold $K$ remains the same for any other occasion. ADVICE deliberately only considers one standard deviation as two or more standard deviations would make the adaptation slow to respond to emerging safety risks. Our sensitivity analysis on these two parameters (Appendix E), corroborates the described behaviours.
@@ 5 Theoretical Analysis
Through theoretical analysis, we show that the misclassification probability of an unseen feature is bounded and can be reduced by diversifying data collected before episode $E$. We also derive regret bounds for ADVICE  with fixed and adaptive $K$, evidencing that long-term regret is lower with an adaptive $K$. All theorem proofs are in Appendix B.
@@
Theorem 1 shows that ADVICE’s misclassification probability decreases exponentially with a larger contrastive separation margin $\gamma$ and smaller noise $\sigma^{2}$, emphasizing the importance of training an efficient encoder for the task.
Theorem 2.
Theorem 2 extends Theorem 1 by highlighting that diverse and representative data during the collection stage reduces misclassification rates, with dependence on the entropy $\sqrt{H(F_{E})}$ of the feature set.
Theorem 3.
The expected regret of ADVICE with a fixed $K$ is bounded by $O\left(\frac{1}{\sqrt{E_{max}}}\right)+O\left(E_{max}\cdot\Delta Q\cdot exp\left(-\frac{\gamma^{2}}{2\sigma^{2}}\right)\cdot\frac{1}{\sqrt{K}}\right)$.
Theorem 4.
The expected long-term regret of ADVICE with an adaptive $K$ is lower than ADVICE that of a fixed $K$ and is bounded by $O\left(\frac{1}{\sqrt{E_{max}}}\right)+O\left(E_{max}\cdot\Delta Q\cdot exp\left(-\frac{\gamma^{2}}{2\sigma^{2}}\right)\cdot\frac{1}{\sqrt{K^{*}}}\right)+O\left(\frac{1}{E_{max}}\sum^{E_{max}}_{e=1}|K_{e}-K^{*}|\cdot\Delta Q_{K_{e}}\right)$.
6 Evaluation
We evaluate ADVICE  [^1] The ADVICE code is available at: https://anonymous.4open.science/r/ADVICE-6AF9 using tasks from the Safety Gymnasium test-suite (Ji et al., 2023b), where a robot with Lidar sensors navigates obstacle-filled environments. This benchmark assesses performance and safety in high-dimensional scenarios with various constraints, enabling a fair comparison of ADVICE against techniques that also used this benchmark (Bharadhwaj et al., 2020). A terminal state occurs when the agent collides with an obstacle, causing catastrophic damage and ending the episode. The selected tasks cover varying complexities in agent, goal, and obstacle positions:
  • Semi-random Goal: A standard goal environment with six static obstacles, with the agent and goal spawning randomly in each episode.
  • Randomised Goal: A similar goal environment, with obstacles also having random positions per episode.
  • Randomised Button: The agent must press the correct button in an environment with six obstacles and four buttons, with random positions each episode. Pressing the wrong button terminates the episode.
  • Constrained Randomised Goal: A randomised goal environment with hazards that impose a step-wise cost when the agent is inside. The aim is to minimise the cost.
@@ 6 Evaluation
Figure 2: Average episodic reward, cumulative safety violations, cumulative goal reaches of examined methods (DDPG, DDPG-Lag, RCPO, Tabular shield, Conservative Safety Critic, ADVICE) and example latent space visualisation for the semi-random goal (a), randomised goal (b), randomised button (c), and randomised circle (d) environments. Results are shown for episodes 10001000100010002000200020002000 to highlight ADVICE’s execution and adaptation (Figure 1-B).
Details of these environments and tasks are in Appendix A, with Figure 6 showing task snapshots from the Safety Gymnasium. To evaluate ADVICE, we compare it with the following state-of-the-art algorithms:
  • DDPG: A deep deterministic policy gradient agent  (Lillicrap et al., 2015) which is the foundational baseline.
  • DDPG-Lag: A DDPG agent with an online Lagrangian multiplier ($\lambda\!=\!0.1$, $\alpha\!=\!0.01$) which heavily penalises constant safety violations (Borkar, 2005) dynamically.
  • RCPO: A reward-constrained policy optimization framework, where a Lagrangian balances reward maximization and constraint satisfaction (Tessler et al., 2018).
  • Tabular Shield: A DDPG agent with a tabular shield storing terminal state-action pairs, preventing actions within 0.1 of recorded pairs (Shperberg et al., 2022).
  • Conservative Safety Critic (CSC): A DDPG agent with a conservative safety critic that uses conservative estimates to evaluate safety (Bharadhwaj et al., 2020).
Other black-box safe exploration research discussed in Section 2 was excluded due to substantial data collection requirements or because they represent prior research from assessed methods used in our evaluation. All methods use the same DDPG configuration for a fair comparison, with algorithm details and hyperparameters in Appendix F. Results are averaged over three runs, showing mean scores and confidence intervals (standard error). Unless stated otherwise, the default deployment ADVICE per run is $E\!=\!1000$, $K=4$, $K^{max}=5$, $T=1000$, $h_{d}=25$, and $h_{r}=3$. To evaluate all safe RL methods, we use a common setup with unconstrained, black-box MDPs, which may result in sparse data due to infrequent constraint violations. Figure 8 in Appendix C evaluates all methods in a constrained environment (Constrained Randomised Goal) to align with the original designs of DDPG-Lag, RCPO, and CSC and assess ADVICE’s performance.
6.1 Performance Results
For our experiments, we evaluate the performance of the comparable methods in the semi-random goal, randomised goal, and randomised circle environments. Figure 2 shows average episodic reward, cumulative safety violations, cumulative goal reaches, and a latent space visualisation from a single ADVICE run. Full results are in Appendix C. The results show that ADVICE, despite prioritising safety, achieves competitive rewards across tasks. Both DDPG-Lag and ADVICE exhibit a trade-off between reward maximisation and safety.
@@ 6.1 Performance Results
Figure 3: Example trajectories of the assessed methods on the semi-random goal environment; obstacles are purple, goals are green circles, and the RL agent is the red vehicle.
@@
The primary goal of a safety shield is to enable safe exploration and learning. Across all tasks, ADVICE significantly reduces safety violations, outperforming other methods by a notable margin. While DDPG-Lag also reduces violations, it is less effective than ADVICE, highlighting the strength of its shield. RCPO and CSC reduce safety violations by a small amount but vastly underestimate safety due to sparse data in unconstrained MDPs with complex constraints, where regular neural network-based techniques struggle without a consistent cost signal. This demonstrates the strength of ADVICE’s contrastive learning. The learnt latent space from an ADVICE run (Figure 2-column 4) further validates its ability to distinguish safe and unsafe features. ADVICE constructs its shield during the first $E\!=\!1000$ episodes before executing safe exploration, whereas DDPG-Lag reduces violations from episode $E\!=\!1$, further affirming ADVICE’s effectiveness.
Despite cautious prioritisation, ADVICE achieves similar goal-reaching frequencies as other methods while maintaining safety. In fact, ADVICE completes the task without compromising safety, and the difference in accumulated reward is due to the reward function design and the reduced time the agent has to complete the task using a cautious path. Through additional experiments with an increased number of maximum steps per episode ($T\!=\!2000$ – Figure 13 - Appendix C), we have established that ADVICE’s average episodic return is very similar to the other methods. Hence, increasing the episode’s duration yields improved ADVICE-based results. These results confirm that ADVICE’s contrastive shielding effectively reduces safety violations compared to other methods, with minimal impact on the agent’s overall performance.
Figure 3 shows example trajectories in the semi-random goal environment, supporting our findings. DDPG and Tabular shield prioritise maximising return, navigating through the centre of the object wall, which is risky due to action noise, as seen in Figure 2(a). Cost-based methods learn similar behaviours but maintain a larger gap from objects to account for noise. In contrast, ADVICE takes a cautious path around the object wall, reducing collision risk more than any method. This longer trajectory slightly lowers the reward due to initially moving further from the goal as to avoid the obstacles as much as possible, thus, explaining the lower average episodic reward observed as the reward functions penalise moving away from the goal in anyway.
@@ 6.2 Transfer Learning
Figure 4 highlights ADVICE’s transfer learning and robustness to slight distributional shift capabilities in a semi-random environment. Using a pre-trained shield from a randomized environment, ADVICE significantly reduces safety violations in a new setting with different goal and agent positions. Over 1000 episodes, it reduces violations by over 50%, outperforming RCPO, Tabular Shield, and CSC and showcasing its ability to generalize the shield to similar/non-stationary environments despite subtle differences. This demonstrates ADVICE’s ability to generalize its shield to similar, non-stationary environments. While ensuring safety, ADVICE maintains comparable reward and goal-reaching frequencies, showcasing its robustness and effectiveness in transfer learning tasks.
@@ 7 Conclusion and Future Work
We introduced ADVICE, a post-shielding technique for the safe exploration of RL agents operating in complex black-box environments. ADVICE does not need any prior knowledge and uses a contrastive autoencoder to distinguish between safe and unsafe features efficiently. Our evaluation shows that ADVICE significantly reduces safety violations while maintaining competitive performance against state-of-the-art methods. In future work plan to explore meta-learning techniques (Hospedales et al., 2021), enabling faster activation of ADVICE without sacrificing performance. Quantisation or pruning could reduce computational demands, enhancing applicability in computationally constrained domains. We envisage that ADVICE serves as a foundational step in using neural network-based shielding for safe RL exploration in complex, black-box environments without any prior knowledge.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.

### v3

@@ 4.2 ADVICE Execution and Adaptation
Upon completing the shield construction phase, ADVICE, similar to [34], can be used as a post-shield within the standard RL loop. The constructed shield encourages safe environment exploration, guarding the agent throughout its interaction with the environment by ensuring the execution of safe actions.
The ADVICE execution and adaptation stage (Algorithm 1, the standard RL loop with the ADVICE extension) involves performing the action evaluation and decision-making cycle from lines 123 until the maximum number of training episodes $E_{max}$ is reached in line 1. During each step $t$ of an episode’s execution (lines 316), the RL agent, having observed the current state $s_{t}$, selects an action $a_{t}$ based on its current policy $\pi$ (line 4). Without the ADVICE shield, the agent would proceed with this action regardless of its potential safety implications. Instead, the ADVICE shield is applied in lines 512 to assess whether the agent-selected action is safe – and to replace it with a safe action if not, thus ensuring safe environment exploration. This assessment is performed by invoking the function IsSafe in line 6. To establish whether the action associated with a feature $f$ is safe, this function (lines 2528) uses the contrastive autoencoder $\mathit{CA}$ to extract the latent representation $\hat{f}$ of that feature (line 26), invokes the function GetSafeNeighbours to collect – from the $K_{max}$ latent data points nearest to $\hat{f}$ – those deemed safe into a set $N_{\hat{f}}$ of safe neighbours (line 27). Finally, the assessed action is deemed safe if the cardinality of the set $N_{\hat{f}}$ is at least as high as the safety threshold $K$, and unsafe otherwise (line 28).
The safety threshold $K\in[0,K_{max}]$ determines how many neighbours need to be labelled safe for the encoded feature to be deemed safe. Values of $K$ close to $K_{\mathit{max}}$ make ADVICE more risk-averse, while $K$ values closer to $\lceil K_{\mathit{max}}/2\rceil$ yield ADVICE shields that are more relaxed and favour exploration. A value of $K<\lceil K_{\mathit{max}}/2\rceil$ should not be considered, as this would allow the RL agent to execute actions that are more likely to be unsafe than safe.
@@
At the end of each episode, ADVICE updates its cautiousness level by considering the recent safety performance of the RL agent (line 22). This unique ADVICE characteristic enables moving beyond the static definition of the safety threshold $K$ by automatically adjusting its value based on its recent safe/unsafe behaviour. Accordingly, ADVICE becomes adaptive and allows the agent to explore more when exhibiting safe behaviour for a period of time while being more cautious, thus interfering more when the RL agent behaves increasingly unsafely.
To assess the agent’s performance over time, ADVICE employs a double sliding window, commonly used in the field of anomaly detectionBy comparing the current trend of safety violations against a broader historical view, ADVICE can discern whether the current trend deviates from the pattern seen historically. This analysis informs ADVICE to strengthen or relax its cautiousness level during the execution of an episode. The adaptation function is given by:
$K=\begin{cases}\min(K\!+\!1,K_{max}),&\text{if }\mu_{r}\!\!>\!\!\mu_{d}\!+\!\sigma_{d}\\ \max(K\!-\!1,\lceil K_{max}/2\rceil),&\text{if }\mu_{r}\!\!<\!\!\mu_{d}-\sigma_{d}\\ K,&\text{otherwise}\end{cases}$ (4)
@@
If the $\mu_{r}$ value exceeds the sum of $\mu_{d}+\sigma_{d}$, this signifies that the agent incurs safety violations more often than expected. Consequently, $K$ is automatically incremented to adopt a more cautious position. Conversely, if $\mu_{r}$ is less than $\mu_{d}-\sigma_{d}$, it is an indication that the agent acts more cautiously than expected. As a result, $K$ is automatically reduced to allow the agent to explore more freely. The safety threshold $K$ remains the same for any other scenario. ADVICE deliberately only considers one standard deviation as two or more standard deviations would make the adaptation slow to respond to emerging safety risks. Our sensitivity analysis on these two parameters (Appendix E), corroborates the described behaviours.
@@ 5 Theoretical Analysis
Through theoretical analysis, we show that the misclassification probability of unseen features is bounded and decreases with more diverse data collected up to episode $E$. We also derive regret bounds for ADVICE with fixed and adaptive $K$, showing that adaptive $K$ yields lower long-term regret. All theorem proofs are in Appendix B.
@@
Figure 2: Average episodic reward, cumulative safety violations, cumulative goal reaches of examined approaches (DDPG, DDPG-Lag, RCPO, Tabular shield, CSC, ADVICE) and example latent visualisation for the semi-random goal (a), randomised goal (b), randomised button (c), and randomised circle (d) environments. Results are shown for episodes 1000100020002000 to highlight ADVICE’s execution and adaptation (Figure 1-B), with full results shown in Appendix C (Figure 9).
Theorem 1 shows that ADVICE’s misclassification probability decreases exponentially with a larger contrastive separation margin $\gamma$ and smaller noise $\sigma^{2}$. This emphasizes the importance of training an efficient encoder during ADVICE’s shield construction stage.
Theorem 2.
Theorem 2 extends Theorem 1 by highlighting that diverse and representative data collected up to episode $E$ reduces misclassification rates, with dependence on the entropy $\sqrt{H(F_{E})}$ of the feature set.
Theorem 3.
The expected regret of ADVICE with a fixed $K$ is upper bounded by a function of order $O\left(\frac{1}{\sqrt{E_{max}}}\right)+O\left(E_{max}\cdot\Delta Q\cdot exp\left(-\frac{\gamma^{2}}{2\sigma^{2}}\right)\cdot\frac{1}{\sqrt{K}}\right)$.
Theorem 4.
The expected long-term regret of ADVICE with an adaptive $K$ is lower than that of ADVICE with a fixed $K$, and is upper bounded by a function of order $O\left(\frac{1}{\sqrt{E_{max}}}\right)+O\left(E_{max}\cdot\Delta Q\cdot exp\left(-\frac{\gamma^{2}}{2\sigma^{2}}\right)\cdot\frac{1}{\sqrt{K^{*}}}\right)+O\left(\frac{1}{E_{max}}\sum^{E_{max}}_{e=1}|K_{e}-K^{*}|\cdot\Delta Q_{K_{e}}\right)$.
@@
Figure 3: Example trajectories of the assessed approaches on the semi-random goal environment; obstacles are purple, goals are green circles, and the RL agent is the red vehicle.
Figure 4: Average episodic reward, cumulative safety violations, and cumulative goal reaches of various approaches on the randomised goal environment where the maximum episodic steps are doubled.
6 Evaluation
We evaluate ADVICE using tasks from the Safety Gymnasium test-suite [25], where a robot with Lidar sensors navigates environments with multiple obstacles. This benchmark assesses performance and safety in high-dimensional scenarios with various constraints, enabling a fair comparison of ADVICE against approaches that also used this benchmark [13]. A terminal state occurs when the agent collides with an obstacle, causing catastrophic damage and ending the episode. The selected tasks cover varying complexities in agent, goal, and obstacle positions:
  • Semi-random Goal: A goal environment with six static obstacles, with the agent and goal spawning randomly in each episode.
  • Randomised Goal: Similar to the Semi-random Goal environment, with obstacles also having random positions per episode.
  • Randomised Button: The agent must press the correct button in an environment with six obstacles and four buttons, with random positions per episode. Pressing the wrong button ends the episode.
  • Constrained Randomised Goal: A randomised goal environment with hazards that impose a step-wise cost when the agent is inside. The aim in this environment is to minimise the cost.
Details of these environments and tasks are in Appendix A, with Figure 8 showing task instances from the Safety Gymnasium.
To evaluate ADVICE, we compare it with the following state-of-the-art algorithms:
  • DDPG: A deep deterministic policy gradient agent  [29] which is the foundational baseline.
  • DDPG-Lag: A DDPG agent with an online Lagrangian multiplier ($\lambda\!=\!0.1$, $\alpha\!=\!0.01$) which heavily penalises constant safety violations [14] dynamically.
  • RCPO: A reward-constrained policy optimisation framework, where a Lagrangian balances reward maximisation and constraint satisfaction [39].
  • Tabular Shield: A DDPG agent with a tabular shield storing terminal state-action pairs, preventing actions within 0.1 of recorded pairs [34].
  • Conservative Safety Critic (CSC): A DDPG agent with a conservative safety critic that uses conservative estimates to evaluate safety [13].
Other black-box safe exploration approaches discussed in Section 2 were excluded because they either represent earlier research from approaches used in our evaluation or due to substantial data collection requirements. To enable a fair comparison against the evaluated approaches, they use the same DDPG configuration; algorithm details and hyperparameter information are provided in Appendix F. Following [16, 13], we perform five independent runs, with results showing mean scores and confidence intervals (standard error). Unless stated otherwise, the default deployment ADVICE per run is E=1000E\!=\!1000, $K=4$, $K_{max}=5$, $T=1000$, $h_{d}=25$, and $h_{r}=3$.
To evaluate safe RL approaches under a common setting, we adopt unconstrained, black-box MDPs, where constraint violations (non-accepting terminal states) are sparse. In this setup, ADVICE collects features upon encountering a non-accepting terminal state, while constrained approaches (e.g., DDPG-Lag, RCPO, CSC) receive explicit cost signals at each violation. Upon encountering a non-accepting terminal state, the episode terminates and is reset. To further ensure fairness, we also evaluate all approaches in a fully constrained MDP environment (Constrained Randomised Goal, see Figure 10), where explicit costs are assigned during an episode. In this setting, ADVICE collects features upon observing a cost, conforming to the assumptions of constrained approaches during training.
6.1 Performance Results
@@
For our experiments, we evaluate the performance of the comparable approaches in all environments. Figure 2 shows average episodic reward, cumulative safety violations, and cumulative goal reaches; the rightmost figure shows the latent space visualisation from a single ADVICE run. Full results are in Appendix C. The results show that ADVICE prioritises safety and achieves competitive rewards across tasks. Both DDPG-Lag and ADVICE exhibit a trade-off between reward maximisation and safety.
The primary goal of a safety shield is to enable safe exploration and learning. Across all tasks, ADVICE significantly reduces safety violations, outperforming other approaches by a notable margin. While DDPG-Lag also reduces violations, it is less effective than ADVICE, highlighting the strength of ADVICE’s shield. RCPO and CSC reduce safety violations by a small amount but vastly underestimate safety due to sparse data in unconstrained MDPs with complex constraints, where regular neural network-based approaches struggle without a consistent cost signal. This demonstrates the strength of ADVICE’s contrastive learning. The learnt latent space from an ADVICE run (Figure 2-column 4) further validates its ability to distinguish safe and unsafe features. Despite requiring an initial unshielded exploration phase (E=1000E\!=\!1000 episodes) to construct its shield, ADVICE ultimately achieves fewer safety violations than DDPG-Lag, which applies its constraints from episode E=1E\!=\!1. This highlights the superior long-term effectiveness of ADVICE’s shielding mechanism.
Despite cautious prioritisation, ADVICE achieves similar goal-reaching frequencies as other approaches while maintaining safety. In fact, ADVICE completes the task without compromising safety, and the difference in accumulated reward is due to the reward function design and the reduced time the agent has to complete the task using a cautious path. Through additional experiments with an increased number of maximum steps per episode (T=2000T\!=\!2000 – Figure 4), we have established that ADVICE’s average episodic return is very similar to the other approaches. Hence, increasing the episode’s duration yields improved ADVICE-based results. These results confirm that ADVICE’s contrastive shielding effectively reduces safety violations compared to other approaches, with minimal impact on the agent’s overall performance.
Figure 3 shows example trajectories in the semi-random goal environment, further supporting our findings. DDPG and Tabular shield prioritise return maximisation, navigating through the centre of the object wall, which is risky due to action noise, as seen in Figure 2a. Cost-based approaches learn similar behaviours but maintain a larger gap from objects to account for noise. In contrast, ADVICE follows a cautious path around the wall, reducing collision risk more than any other method. This longer trajectory lowers the average episodic reward, since the agent initially moves away from the goal to avoid obstacles, and the reward function penalises such detours while neglecting safety risk.
@@ 6.2 Transfer Learning
Figure 5 shows ADVICE’s transfer learning and robustness under distribution shift in a semi-random environment. Using a pre-trained shield from the randomized setting, it cuts safety violations by over 50% across 1000 episodes, clearly outperforming RCPO, Tabular Shield, and CSC. These results demonstrate ADVICE’s ability to generalise its learned shield to similar, non-stationary environments while maintaining comparable reward and goal-reaching rates, highlighting both robustness and effectiveness.
6.3 Constrained Environment
We also demonstrate ADVICE’s ability to operate in unconstrained MDP environments (albeit being less common than constrained ones). Figure 6 shows results in the Constrained Randomised Goal environment (full results in Appendix C). ADVICE consistently achieved the lowest average episodic cost, outperforming cost-based approaches, which showed significant oscillations, highlighting their difficulty in balancing cost reduction with reward maximization [30]. Notably, such oscillations were absent in sparse-cost environments like the unconstrained MDP environments discussed in Section 6.1.
@@ 7 Conclusion and Future Work
ADVICE is a post-shielding approach for the safe exploration of RL agents operating in complex black-box environments. ADVICE does not need any domain knowledge or any scenario-specific safety definitions, thus yielding improved generalizability. Our evaluation shows that ADVICE significantly reduces safety violations while maintaining competitive performance against state-of-the-art approaches. Our future work involves exploring meta-learning techniques for faster ADVICE activation without sacrificing performance. Quantisation could reduce computational demands, enhancing applicability in computationally constrained domains, while using uncertainty quantification [1, 12] in state estimation could yield more robust policies. ADVICE could become a foundational step in using neural network-based shielding for safe RL exploration in complex, black-box environments without any domain knowledge.
{ack}
This research has received funding from the Doctoral Centre for Safe, Ethical and Secure Computing (SEtS) at the University of York, UK, the European Union’s Horizon projects GuardAI and AI4Work (grant agreements No 101168067 and 101135990, respectively), and the UK Advanced Research and Invention Agency’s Safeguarded AI project ULTIMATE. This work was also supported by the Centre for Assuring Autonomy, University of York, UK. The authors are grateful to Charmaine Barker for her help and insight on this paper.